# Single-loop Algorithms for Stochastic Non-convex Optimization with Weakly-Convex Constraints

### v1

@@ 1 Introduction
In this paper, we consider solving the following non-linear inequality constrained optimization problem:
@@
where $F$ and $h_{k}$, $k=1,\dots,m$ are stochastic, weakly convex and potentially non-smooth. This general formulation captures a wide range of practical optimization problems where non-smoothness arises in the objective and constraint functions. Inequality constrained optimization arise in many applications in ML/AI, including continual learning [21], fairness-aware learning [35], multi-class Neyman-Pearson classification (mNPC) problem [27], robust learning [30], policy optimization problem for reinforcement learning [33], and distributed data center scheduling problem [39].
In optimization literature, many techniques have been studied for solving convex constrained optimization problems, including primal-dual methods [6, 29], augmented Lagrangian methods [5], penalty methods [7], and level-set methods [24]. However, the presence of non-convexity in the problem often complicates algorithmic design and convergence analysis, while non-smoothness and stochasticity add additional layers of difficulty.
Several recent studies have explored stochastic non-convex constrained optimization under various assumptions. Ma et al. [27] and Boob et al. [6] initiated the study of stochastic optimization with weakly convex and non-smooth objective and constraint functions. Inspired by proximal-point methods for weakly-convex unconstrained optimization [10], they have proposed a double-loop algorithm, where a sequence of convex quadratically regularized subproblems is solved approximately. They have derived the state-of-the-art complexity in the order of $O(1/\epsilon^{6})$ for finding a nearly $\epsilon$-KKT solution under a uniform Slater’s condition and a strong feasibility condition, respectively.
Recently, Alacaoglu and Wright [1], Li et al. [21], Huang and Lin [17] have proposed single-loop algorithms for stochastic non-convex constrained optimization. However, these results either are restricted to smooth problems with equality constraints [1] or suffer from an iteration complexity higher than $O(1/\epsilon^{6})$ [21, 17]. This raises an intriguing question: “Can we design a single-loop method for solving non-smooth weakly-convex stochastic constrained optimization problems to match the state-of-the-art complexity of $O(1/\epsilon^{6})$ for finding a nearly $\epsilon$-KKT solution?"
Table 1: Comparison with existing works. We consider that both objective function and constraint function oracles are stochastic. In the Constraint column, $h_{k}$ means multiple constraints, $k$ from 1 to $m$. In Convexity column, NC means Non-convex, WC means Weakly Convex, C means Convex. The monotonicity column means non-decreasing. The last column Complexity means the complexity for finding an (nearly) $\epsilon$-KKT point of (1).
@@
[1] Single Loop $F(\cdot)$ $h(\cdot)=0$ $F,h$ NC ($F,h$) none $\tilde{\mathcal{O}}(\epsilon^{-5})$
[23] Double Loop $F(\cdot)+g(\cdot)$ $h(\cdot)=0$ $F,h$ C ($g$), NC ($F,h$) none $\tilde{\mathcal{O}}(\epsilon^{-5})$
[27] Double Loop $F(\cdot)$ $h_{k}(\cdot)\leq 0$ none WC ($F,h_{k}$) none $\mathcal{O}(\epsilon^{-6})$
[6] Double Loop $F(\cdot)$ $h_{k}(\cdot)\leq 0$ none WC ($F,h_{k}$) none $\mathcal{O}(\epsilon^{-6})$
[17] Single Loop $F(\cdot)$ $h(\cdot)\leq 0$ none WC ($F$), C ($h$) none $\tilde{\mathcal{O}}(\epsilon^{-8})$
[21] Single Loop $\sum_{i}f_{i}(g_{i}(\cdot))$ $h_{k}(\cdot)\leq 0$ $f_{i},g_{i},h_{k}$ NC ($f_{i},g_{i},h_{k}$) none $\mathcal{O}(\epsilon^{-7})$
@@
We answer the question in the affirmative. We propose a novel algorithm based on the well-studied exact penalty method [42, 13]. Different from Alacaoglu and Wright [1], Li et al. [21] who used a quadratic or squared hinge penalty, we construct a penalty of the constraints using the non-smooth hinge function, i.e., by adding $\frac{1}{m}\sum_{k=1}^{m}\beta[h_{k}(\mathbf{x})]_{+}$ into the objective, where $\beta$ is a penalty parameter. Despite being non-smooth, the benefits of this penalty function include (i) it preserves the weak convexity of in the penalty; (ii) it allows using a constant $\beta$ as opposed to a large $\beta$ required by the squared hinge penalty [21]; (iii) it introduces the structure of finite-sum coupled compositional optimization (FCCO), enabling the use of existing FCCO techniques in the algorithmic design. Consequently, this approach establishes state-of-the-art complexity for achieving a nearly $\epsilon$-KKT solution, under a regularity assumption similar to those in [21, 1]. We summarize our contributions as follows.
@@ 2 Related Work
Despite their long history of study, recent literature on optimization with non-convex constraints has introduced new techniques and theories for the penalty method [25], Lagrangian methods [31, 34, 23, 22], the sequential quadratic programming methods [3, 4, 8, 28] and trust region method [14]. These works mainly focused on the problems with smooth objective and constraint functions. On the contrary, based on Moreau envelopes, Ma et al. [27], Boob et al. [6], Huang and Lin [17], Jia and Grimmer [19] have developed algorithms and theories for problems with non-smooth objective and constraint functions. However, these methods cannot be applied when the objective function has a compositional structure. The method by Li et al. [21] can be applied when the objective function has a coupled compositional structure, but it requires the objective and constraint functions to be smooth. Compared to their work, we relax the smoothness assumption, but assume the monotonicity of the outer function in the compositional structure.
@@
Alacaoglu and Wright [1] studies a single-loop quadratic penalty method for stochastic smooth nonconvex equality constrained optimization under a generalized full-rank assumption on the Jacobian of the constraints. Under a similar assumption, Li et al. [23] proposed a double-loop inexact augmented Lagrangian method using the momentum-based variance-reduced proximal stochastic gradient algorithm as a subroutine. Both Alacaoglu and Wright [1] and Li et al. [23] consider only smooth equality constraints and both achieved a complexity of $\mathcal{O}(1/\epsilon^{5})$.  [^1] Complexity lower than $\mathcal{O}(1/\epsilon^{5})$ can be achieved by Alacaoglu and Wright [1] when the constraints are deterministic. Ma et al. [27] and Boob et al. [6] studied weakly convex non-smooth optimization with inequality constraints. They proposed double-loop methods, where a sequence of proximal-point subproblems is solved approximately by a stochastic subgradient method [39] or a constraint extrapolation method [6]. Under a uniform Slater’s condition and a strong feasibility assumption, respectively, Ma et al. [27] and Boob et al. [6] establish a complexity of $\mathcal{O}(1/\epsilon^{6})$. Under the uniform Slater’s condition, Huang and Lin [17] proposed a single-loop switching subgradient method for non-smooth weakly convex constrained optimization problems, but the complexity of their method is $\mathcal{O}(1/\epsilon^{8})$ when the constraints are stochastic.
@@
Motivated by applications from continual learning with non-forgetting constraints, Li et al. [21] introduced a single-loop squared hinge penalty method for a FCCO [36] problem, where the objective function takes the form of $\sum_{i}f_{i}(g_{i}(\cdot))$ and the constraints are non-convex and smooth. They establish a complexity of $\mathcal{O}(1/\epsilon^{7})$ under the generalized full-rank assumption on the Jacobian of constraints. In our work, we consider a FCCO problem with nonconvex but unnecessarily smooth inequality constraints. Following the traditional exact penalty method, we first add a hinge penalty of the constraints to the objective function, and solve the unconstrained FCCO using a single-loop algorithm. We show that our method improves the complexity to $\mathcal{O}(1/\epsilon^{6})$ in two cases: (1) $f_{i}$, $g_{i}$ and $h_{k}$ are non-smooth weakly convex and $f_{i}$ is monotone non-decreasing, and (2) $f_{i}$ and $g_{i}$ are smooth nonconvex but $h_{k}$ is non-smooth weakly convex. Note that the second case includes the special case where $f_{i}$, $g_{i}$ and $h_{k}$ are smooth non-convex. A comparison with prior results in presented in Table 1.
@@
Finally, we would like to acknowledge a concurrent work by [26]  [^2] The first version of the present work [37] was submitted on 01/30/2025.. Their study also explores a penalty method for problems with weakly convex objective and constraint function. The key differences between our work and theirs are: (i) they consider only one constraint function, while we tackle multiple constraint functions with stochastic sampling; (ii) they do not directly use the hinge penalty function, but instead adopt a smoothed version of it; and (iii) they employ the SPIDER/SARAH technique to track the constraint function value, which requires periodically using large batch sizes. In contrast, we track multiple constraint function values by employing the MSVR technique [20] using only a constant batch size.
@@ 3 Notations and Preliminaries
[…]
We say $\psi$ is $\mu$-strongly convex ($\mu\geq 0$) if $\psi(\mathbf{x})\geq\psi(\mathbf{x}^{\prime})+\langle\zeta,\mathbf{x}-\mathbf{x}^{\prime}\rangle+\frac{\mu}{2}\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}$ for any $(\mathbf{x},\mathbf{x}^{\prime})$ and any $\zeta\in\partial\psi(\mathbf{x}^{\prime})$. We say $\psi$ is $\rho$-weakly convex ($\rho>0$) if $\psi(\mathbf{x})\geq\psi(\mathbf{x}^{\prime})+\langle\zeta,\mathbf{x}-\mathbf{x}^{\prime}\rangle-\frac{\rho}{2}\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}$ for any $(\mathbf{x},\mathbf{x}^{\prime})$ and any $\zeta\in\partial\psi(\mathbf{x}^{\prime}).$ We say $\psi$ is $L$-smooth ($L\geq 0$) if $|\psi(\mathbf{x})-\psi(\mathbf{x}^{\prime})+\langle\nabla\psi(\mathbf{x}^{\prime}),\mathbf{x}-\mathbf{x}^{\prime}\rangle|\leq\frac{L}{2}\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}$ for any $(\mathbf{x},\mathbf{x}^{\prime})$. For simplicity, we abuse $\partial\psi(x)$ to denote one subgradient from the corresponding subgradient set when no confusion could be caused.
@@
For a non-convex optimization problem, finding a globally optimal solution is intractable. Instead, a Karush-Kuhn-Tucker (KKT) solution is of interest, which is an extension of a stationary solution of an unconstrained non-convex optimization problem. A solution $\mathbf{x}$ is a KKT solution to (1) if there exist $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{m})^{\top}\in\mathbb{R}_{+}^{m}$ such that $0\in\partial F(\mathbf{x})+\sum_{k=1}^{m}\lambda_{k}\partial h(\mathbf{x})$, $h_{k}(\mathbf{x})\leq 0,\forall k$ and $\lambda_{k}h_{k}(\mathbf{x})=0$ $\forall k$. For non-asymptotic analysis, one might consider finding an $\epsilon$-KKT solution as defined below.
@@
A solution $\mathbf{x}$ is an $\epsilon$-KKT solution to (1) if there exist $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{m})^{\top}\in\mathbb{R}_{+}^{m}$ such that $\text{dist}(0,\partial F(\mathbf{x})+\sum_{k=1}^{m}\lambda_{k}\partial h(\mathbf{x}))\leq\epsilon$, $h_{k}(\mathbf{x})\leq\epsilon,\forall k$, and $\lambda_{k}h_{k}(\mathbf{x})\leq\epsilon$ if $h_{k}(\mathbf{x})\geq 0$ or $\lambda_{k}=0$ if $h_{k}(\mathbf{x})<0$ for $k=1,\dots,m$.
However, since the objective and the constraint functions are non-smooth, finding an $\epsilon$-KKT solution is not tractable, even the constraint functions are absent. For example, if $F(x)=|x|$ finding $\epsilon$-stationary solution is infeasible unless we find the optimal solution $x=0$. To address this challenge, we follow many previous works on weakly convex optimization [10] and consider finding a nearly $\epsilon$-KKT solution defined below.
@@
A solution $\mathbf{x}$ is a nearly $\epsilon$-KKT solution to (1) if there exist $\bar{\mathbf{x}}$ and $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{m})^{\top}\in\mathbb{R}_{+}^{m}$ such that (i) $\|\mathbf{x}-\bar{\mathbf{x}}\|\leq O(\epsilon)$, $\text{dist}(0,\partial F(\bar{\mathbf{x}})+\sum_{k=1}^{m}\lambda_{k}\partial h(\bar{\mathbf{x}}))\leq\epsilon$, (ii) $h_{k}(\bar{\mathbf{x}})\leq\epsilon,\forall k$, and (iii) $\lambda_{k}h_{k}(\bar{\mathbf{x}})\leq\epsilon$ if $h_{k}(\bar{\mathbf{x}})\geq 0$ or $\lambda_{k}=0$ if $h_{k}(\bar{\mathbf{x}})<0$ for $k=1,\dots,m$.
In order to develop our analysis, we introduce the Moreau envelope of a $\rho$-weakly convex function $\phi$:
$\phi_{\theta}(\mathbf{x}):=\min_{\mathbf{y}}\left\{\phi(\mathbf{y})+\frac{1}{2\theta}\|\mathbf{y}-\mathbf{x}\|^{2}\right\}$
where $\theta<\rho^{-1}$.
The Moreau envelope is an implicit smoothing of the original problem. The following lemma follows from standard result [9].
Lemma 3.4.
Hence, if we find $\mathbf{x}$ such that $\|\nabla\phi_{\theta}(\mathbf{x})\|\leq\epsilon$, then we can say that $\mathbf{x}$ is close to a point $\bar{\mathbf{x}}$ that is $\epsilon$-stationary. We call $\mathbf{x}$ a nearly $\epsilon$-stationary solution of $\min_{\mathbf{x}}\phi(\mathbf{x})$.

### v2

@@ 1 Introduction
This paper focuses on solving the following non-linear inequality constrained optimization problem:
@@
where $F$ and $h_{k}$, $k=1,\dots,m$ are stochastic, weakly convex and potentially non-smooth. This general formulation captures a wide range of practical optimization problems where non-smoothness arises in the objective and constraint functions. Inequality constrained optimization arise in many applications in ML/AI, including continual learning (Li et al., 2024a), fairness-aware learning (Vogel et al., 2021), multi-class Neyman-Pearson classification (mNPC) problem (Ma et al., 2020), robust learning (Robey et al., 2021), policy optimization problem for reinforcement learning (Schulman et al., 2015), and distributed data center scheduling problem (Yu et al., 2017).
In optimization literature, many techniques have been studied for solving convex constrained optimization problems, including primal-dual methods (Huang et al., 2025; Boob et al., 2023; Pu et al., 2024; Nedić and Ozdaglar, 2009), augmented Lagrangian methods (Birgin and Martnez, 2014), penalty methods (Boyd and Vandenberghe, 2004), and level-set methods (Lin et al., 2018). However, the presence of non-convexity in the problem often complicates algorithmic design and convergence analysis, while non-smoothness and stochasticity add additional layers of difficulty.
Several recent studies have explored stochastic non-convex constrained optimization under various assumptions. Ma et al. (2020) and Boob et al. (2023) initiated the study of stochastic optimization with weakly convex and non-smooth objective and constraint functions. Inspired by proximal-point methods for weakly-convex unconstrained optimization (Davis and Grimmer, 2019), they have proposed a double-loop algorithm, where a sequence of convex quadratically regularized subproblems is solved approximately. They have derived the state-of-the-art complexity in the order of $O(1/\epsilon^{6})$ for finding a nearly $\epsilon$-KKT solution under a uniform Slater’s condition and a strong feasibility condition, respectively.
Recently, Alacaoglu and Wright (2024); Li et al. (2024a); Huang and Lin (2023) have proposed single-loop algorithms for stochastic non-convex constrained optimization. However, these results either are restricted to smooth problems with equality constraints (Alacaoglu and Wright, 2024) or suffer from an iteration complexity higher than $O(1/\epsilon^{6})$ (Li et al., 2024a; Huang and Lin, 2023). This raises an intriguing question: “Can we design a single-loop method for solving non-smooth weakly-convex stochastic constrained optimization problems to match the state-of-the-art complexity of $O(1/\epsilon^{6})$ for finding a nearly $\epsilon$-KKT solution?"
Table 1: Comparison with existing works. We consider that both objective function and constraint function oracles are stochastic. In the Constraint column, $h_{k}$ means multiple constraints, $k$ from 1 to $m$. In Convexity column, NC means Non-convex, WC means Weakly Convex, C means Convex. The monotonicity column means non-decreasing. The last column Complexity means the complexity for finding an (nearly) $\epsilon$-KKT point of equation 1.
@@
(Alacaoglu and Wright, 2024) Single Loop $F(\cdot)$ $h(\cdot)=0$ $F,h$ NC ($F,h$) none $\tilde{\mathcal{O}}(\epsilon^{-5})$
(Li et al., 2024b) Double Loop $F(\cdot)+g(\cdot)$ $h(\cdot)=0$ $F,h$ C ($g$), NC ($F,h$) none $\tilde{\mathcal{O}}(\epsilon^{-5})$
(Ma et al., 2020) Double Loop $F(\cdot)$ $h_{k}(\cdot)\leq 0$ none WC ($F,h_{k}$) none $\mathcal{O}(\epsilon^{-6})$
(Boob et al., 2023) Double Loop $F(\cdot)$ $h_{k}(\cdot)\leq 0$ none WC ($F,h_{k}$) none $\mathcal{O}(\epsilon^{-6})$
(Huang and Lin, 2023) Single Loop $F(\cdot)$ $h(\cdot)\leq 0$ none WC ($F$), C ($h$) none $\tilde{\mathcal{O}}(\epsilon^{-8})$
(Li et al., 2024a) Single Loop $\sum_{i}f_{i}(g_{i}(\cdot))$ $h_{k}(\cdot)\leq 0$ $f_{i},g_{i},h_{k}$ NC ($f_{i},g_{i},h_{k}$) none $\mathcal{O}(\epsilon^{-7})$
(Liu and Xu, 2025) Single Loop $F(\cdot)$ $h_{k}(\cdot)\leq 0$ none WC ($F,h_{k}$) none $\mathcal{O}(\epsilon^{-6})$
@@
We answer the question in the affirmative. We propose a novel algorithm based on the well-studied exact penalty method (Zangwill, 1967; Evans et al., 1973). Different from Alacaoglu and Wright (2024); Li et al. (2024a) who used a quadratic or squared hinge penalty, we construct a penalty of the constraints using the non-smooth hinge function, i.e., by adding $\frac{1}{m}\sum_{k=1}^{m}\beta[h_{k}(\mathbf{x})]_{+}$ into the objective, where $\beta$ is a penalty parameter. Despite being non-smooth, the benefits of this penalty function include (i) it preserves the weak convexity of in the penalty; (ii) it allows using a constant $\beta$ as opposed to a large $\beta$ required by the squared hinge penalty (Li et al., 2024a); (iii) it introduces the structure of finite-sum coupled compositional optimization (FCCO), enabling the use of existing FCCO techniques in the algorithmic design. Consequently, this approach establishes state-of-the-art complexity for achieving a nearly $\epsilon$-KKT solution, under a regularity assumption similar to those in Li et al. (2024a); Alacaoglu and Wright (2024). We summarize our contributions as follows.
@@ 2 Related Work
Despite their long history of study, recent literature on optimization with non-convex constraints has introduced new techniques and theories for the penalty method (Lin et al., 2022), Lagrangian methods (Sahin et al., 2019; Sun and Sun, 2024; Li et al., 2024b; 2021), the sequential quadratic programming methods (Berahas et al., 2021; 2025; Curtis et al., 2024; Na et al., 2023) and trust region method (Fang et al., 2024). These works mainly focused on the problems with smooth objective and constraint functions. On the contrary, based on Moreau envelopes, Ma et al. (2020); Boob et al. (2023); Huang and Lin (2023); Jia and Grimmer (2022) have developed algorithms and theories for problems with non-smooth objective and constraint functions. However, these methods cannot be applied when the objective function has a compositional structure. The method by Li et al. (2024a) can be applied when the objective function has a coupled compositional structure, but it requires the objective and constraint functions to be smooth. Compared to their work, we relax the smoothness assumption, but assume the monotonicity of the outer function in the compositional structure.
@@
Alacaoglu and Wright (2024) studies a single-loop quadratic penalty method for stochastic smooth nonconvex equality constrained optimization under a generalized full-rank assumption on the Jacobian of the constraints. Under a similar assumption, Li et al. (2024b) proposed a double-loop inexact augmented Lagrangian method using the momentum-based variance-reduced proximal stochastic gradient algorithm as a subroutine. Both Alacaoglu and Wright (2024) and Li et al. (2024b) consider only smooth equality constraints and hence are not applicable to non-smooth constraint functions. Ma et al. (2020) and Boob et al. (2023) studied weakly convex non-smooth optimization with inequality constraints. They proposed double-loop methods, where a sequence of proximal-point subproblems is solved approximately by a stochastic subgradient method (Yu et al., 2017) or a constraint extrapolation method (Boob et al., 2023). Under a uniform Slater’s condition and a strong feasibility assumption, respectively, Ma et al. (2020) and Boob et al. (2023) establish a complexity of $\mathcal{O}(1/\epsilon^{6})$. Under the uniform Slater’s condition, Huang and Lin (2023) proposed a single-loop switching subgradient method for non-smooth weakly convex constrained optimization problems, but the complexity of their method is $\mathcal{O}(1/\epsilon^{8})$ when the constraints are stochastic.
@@
Motivated by applications from continual learning with non-forgetting constraints, Li et al. (2024a) introduced a single-loop squared hinge penalty method for a FCCO (Wang and Yang, 2022) problem, where the objective function takes the form of $\sum_{i}f_{i}(g_{i}(\cdot))$ and the constraints are non-convex and smooth. They establish a complexity of $\mathcal{O}(1/\epsilon^{7})$ under the generalized full-rank assumption on the Jacobian of constraints. Our results can be easily extended to a FCCO problem with nonconvex but unnecessarily smooth inequality constraints. We show that our method improves the complexity to $\mathcal{O}(1/\epsilon^{6})$ in two cases: (1) $f_{i}$, $g_{i}$ and $h_{k}$ are non-smooth weakly convex and $f_{i}$ is monotone non-decreasing, and (2) $f_{i}$ and $g_{i}$ are smooth nonconvex but $h_{k}$ is non-smooth weakly convex. Note that the second case includes the special case where $f_{i}$, $g_{i}$ and $h_{k}$ are smooth non-convex. A comparison with prior results in presented in Table 1.
We also note that some works have considered transferring inequality constraints into equality constraints by adding a slack variable for each constraint (Fukuda and Fukushima, 2017; Ding and Wright, 2023; Li et al., 2021). However, these approaches either require stronger second-order conditions to ensure that a KKT solution to the equality-constrained problem also applies to the original inequality-constrained problem (Fukuda and Fukushima, 2017; Ding and Wright, 2023), or they rely on the boundedness of constraint functions to guarantee this transfer (Li et al., 2021) (cf. Section 4.3 for more discussions). Even when such a transformation is feasible, algorithms for non-smooth stochastic constraint functions remain underdeveloped, underscoring the uniqueness and significance of our contributions.
@@
Finally, we would like to acknowledge a concurrent work by Liu and Xu (2025). Their study also explores a penalty method for problems with weakly convex objective and constraint function. The key differences between our work and theirs are: (i) although they consider a vector-valued function consisting of multiple components, they treat multiple constraint as a single one, i.e., they need to compute mini-batch estimators of all constraint functions at each iteration in their algorithm. In contrast, we do sampling on the constraint functions, and only require estimating $O(1)$ constraint functions at each iteration. (ii) they do not directly use the hinge penalty function, but instead adopt a smoothed version of it; and (iii) they employ the SPIDER/SARAH technique to track the constraint function value, which requires periodically using large batch sizes. In contrast, we track multiple constraint function values by employing the MSVR technique (Jiang et al., 2022) using only a constant batch size.
@@ 3 Notations and Preliminaries
Let $\|\cdot\|$ be the $\ell_{2}$-norm. For $\psi:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$, the subdifferential of $\psi$ at $\mathbf{x}$ is $\partial\psi(\mathbf{x})=\{\zeta\in\mathbb{R}^{d}\mid\psi(\mathbf{x}^{\prime})\geq\psi(\mathbf{x})+\langle\zeta,\mathbf{x}^{\prime}-\mathbf{x}\rangle+o(\|\mathbf{x}^{\prime}-\mathbf{x}\|),\mathbf{x}^{\prime}\to\mathbf{x}\}$, and $\zeta\in\partial\psi(\mathbf{x})$ denotes a subgradient of $\psi$ at $\mathbf{x}$. Let $\cdot]_{+}:=\max\{0,\cdot\}$ denote the hinge function. We say $\psi$ is $\mu$-strongly convex ($\mu\geq 0$) if $\psi(\mathbf{x})\geq\psi(\mathbf{x}^{\prime})+\langle\zeta,\mathbf{x}-\mathbf{x}^{\prime}\rangle+\frac{\mu}{2}\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}$ for any $(\mathbf{x},\mathbf{x}^{\prime})$ and any $\zeta\in\partial\psi(\mathbf{x}^{\prime})$. We say $\psi$ is $\rho$-weakly convex ($\rho>0$) if $\psi(\mathbf{x})\geq\psi(\mathbf{x}^{\prime})+\langle\zeta,\mathbf{x}-\mathbf{x}^{\prime}\rangle-\frac{\rho}{2}\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}$ for any $(\mathbf{x},\mathbf{x}^{\prime})$ and any $\zeta\in\partial\psi(\mathbf{x}^{\prime}).$ We say $\psi$ is $L$-smooth ($L\geq 0$) if $|\psi(\mathbf{x})-\psi(\mathbf{x}^{\prime})+\langle\nabla\psi(\mathbf{x}^{\prime}),\mathbf{x}-\mathbf{x}^{\prime}\rangle|\leq\frac{L}{2}\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}$ for any $(\mathbf{x},\mathbf{x}^{\prime})$. For simplicity, we abuse $\partial\psi(x)$ to denote one subgradient from the corresponding subgradient set when no confusion could be caused.
@@
For a non-convex optimization problem, finding a globally optimal solution is intractable. Instead, a Karush-Kuhn-Tucker (KKT) solution is of interest, which is an extension of a stationary solution of an unconstrained non-convex optimization problem. A solution $\mathbf{x}$ is a KKT solution to (1) if there exist $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{m})^{\top}\in\mathbb{R}_{+}^{m}$ such that $0\in\partial F(\mathbf{x})+\sum_{k=1}^{m}\lambda_{k}\partial h(\mathbf{x})$, $h_{k}(\mathbf{x})\leq 0,\forall k$ and $\lambda_{k}h_{k}(\mathbf{x})=0$ $\forall k$. Of interest in non-asymptotic analysis is finding an $\epsilon$-KKT solution given below.
@@
A solution $\mathbf{x}$ is an $\epsilon$-KKT solution to (1) if there exist $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{m})^{\top}\in\mathbb{R}_{+}^{m}$ such that $\text{dist}(0,\partial F(\mathbf{x})+\sum_{k=1}^{m}\lambda_{k}\partial h(\mathbf{x}))\leq\epsilon$, $h_{k}(\mathbf{x})\leq\epsilon,\forall k$, and $\lambda_{k}h_{k}(\mathbf{x})\leq\epsilon$ if $h_{k}(\mathbf{x})\geq 0$ or $\lambda_{k}=0$ if $h_{k}(\mathbf{x})<0$ for $k=1,\dots,m$.
However, since the objective and the constraint functions are non-smooth, finding an $\epsilon$-KKT solution is not tractable, even the constraint functions are absent. Let us consider a simple example $\min_{\mathbf{x}}|\mathbf{x}|$. The only stationary point is the optimal solution $\mathbf{x}_{*}=0$, and any $\mathbf{x}\neq 0$ is not an $\epsilon$-stationary solution ($\epsilon<1$) no matter how close $\mathbf{x}$ to 0. In other words, unless the iterate generated by the algorithm can exactly land on $\mathbf{x}=0$, it will not produce an $\epsilon$-stationary point. To address this issue, an effective approach for solving non-smooth optimization is to approximate the original problem by a smoothed one using different smoothing techniques, including Nesterov’s smoothing(Nesterov, 2005), randomized smoothing (Kornowski and Shamir, 2022), and Moreau envelope (Davis and Grimmer, 2019). Based on Moreau envelopes, Ma et al. (2020); Boob et al. (2023); Huang and Lin (2023); Jia and Grimmer (2022) have developed algorithms and theories for problems with non-smooth objective and constraint functions, which derives a nearly $\epsilon$-KKT solution. We follow Moreau envelopes technique on non-smooth weakly convex optimization and consider finding a nearly $\epsilon$-KKT solution defined below.
@@
A solution $\mathbf{x}$ is a nearly $\epsilon$-KKT solution to (1) if there exist $\bar{\mathbf{x}}$ and $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{m})^{\top}\in\mathbb{R}_{+}^{m}$ such that (i) $\|\mathbf{x}-\bar{\mathbf{x}}\|\leq O(\epsilon)$, $\text{dist}(0,\partial F(\bar{\mathbf{x}})+\sum_{k=1}^{m}\lambda_{k}\partial h(\bar{\mathbf{x}}))\leq\epsilon$, (ii) $h_{k}(\bar{\mathbf{x}})\leq\epsilon,\forall k$, and (iii) $\lambda_{k}h_{k}(\bar{\mathbf{x}})\leq\epsilon$ if $h_{k}(\bar{\mathbf{x}})\geq 0$ or $\lambda_{k}=0$ if $h_{k}(\bar{\mathbf{x}})<0$ for $k=1,\dots,m$.
In order to develop our analysis, we introduce the Moreau envelope of a $\rho$-weakly convex function $\phi$: $\phi_{\theta}(\mathbf{x}):=\min_{\mathbf{y}}\left\{\phi(\mathbf{y})+\frac{1}{2\theta}\|\mathbf{y}-\mathbf{x}\|^{2}\right\}$, where $\theta<\rho^{-1}$. The Moreau envelope is an implicit smoothing of the original problem (cf. Lemma B.1 in Appendix). Hence, if we find $\mathbf{x}$ such that $\|\nabla\phi_{\theta}(\mathbf{x})\|\leq\epsilon$, then we can say that $\mathbf{x}$ is close to a point $\bar{\mathbf{x}}$ that is $\epsilon$-stationary (Davis and Grimmer, 2019). We call $\mathbf{x}$ a nearly $\epsilon$-stationary solution of $\min_{\mathbf{x}}\phi(\mathbf{x})$.
Remark: Our current formulation focuses on the non-smooth case, there is also a simple extension in the smooth setting. Suppose the objective and constraint functions are smooth, using the Lipschitz contiuity of the gradients, and that of constraint functions, it is easy to prove that a nearly $\epsilon$-stationary point is also an $O(\epsilon)$-stationary point. This is because $\|\mathbf{x}-\bar{\mathbf{x}}\|_{2}\leq\mathcal{O}(\epsilon)$ and leveraging the Lipschitz conunity condition, every condition in the definition of nearly $\epsilon$-KKT solution can be converted to that on $\mathbf{x}$.